# CMSO-transducing tree-like graph decompositions

### v2

@@ 6 Conclusion
We provide transductions for obtaining tree-like graph decompositions such as modular decompositions, cotrees, split decompositions and bi-join decompositions from a graph using $\operatorname{C}{\!\!}\operatorname{MSO}$. This improves upon results of Courcelle [10, 12] who gave such transductions for ordered graphs. It is worth mentionning that Theorem 5.5 can be also used to $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce canonical decompositions of other structures such as Tutte’s decomposition of matroids or generally split-decompositions of submodular functions [16] or modular decompositions of $2$-structures [19] or of hypergraphs [22]. As shown by the application given in [12] for transducing Whitney’s isomorphism class of a graph, a line of research is to more investigate which structures can be $\operatorname{C}{\!\!}\operatorname{MSO}$-transduced from a graph or a set system by using the transductions from Theorem 1.1. Also, naturally, the question arises whether counting is necessary or whether $\operatorname{MSO}$ is sufficient to transduce such decompositions. Furthermore, our results include that transducing rank decompositions for graphs of rank-width 1 is possible using $\operatorname{C}{\!\!}\operatorname{MSO}$. But it is not known whether rank-decompositions can be transduced in general using some suitable extension of $\operatorname{MSO}$. Nevertheless, a corollary of our results is that $\operatorname{C}{\!\!}\operatorname{MSO}$-transducing rank-decompositions can be now reduced to consider $\operatorname{C}{\!\!}\operatorname{MSO}$-transducing rank-decompositions of prime graphs wrt either modular decomposition or split-decomposition as if a graph has rank-width at least $2$, then its rank-width is equal to the rank-width of its prime induced graphs wrt either modular or split decomposition. Thus, our results imply that, for any fixed $k$, there is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction that computes a clique-decomposition of small width for any graph belonging to a graph class whose prime graphs have sizes bounded by $k$ or prime graphs have linear clique-width bounded by $k$, e.g., many $H$-free graphs have small prime graphs or have small linear clique-width (see for instance [6] or [24] for prominent such examples).
Appendix A Deferred proofs of Section 5
Proof A.1 (Proof of Lemma 5.1).
To argue the second assertion, assume that $T$ is the laminar tree of the set system $(U,\mathcal{S}_{a})$. First observe that in the set system $(U,\mathcal{S}_{a})$ the element $a$ is not contained in any set apart from $U$ and $\{a\}$ and hence the leaf $\{a\}$ is a child of the root $U$ of $T$. Additionally, the set $U\setminus\{a\}$ is in $S_{a}$ as the bipartition ($(\{a\},U\setminus\{a\})$ is in $\mathcal{B}$ by definition and hence $U\setminus\{a\}$ is also a child of the root $U$ of $T$. Let $T^{\prime}$ be the tree obtained from $T$ by removing node $U$, making all edges undirected and adding an edge from node $\{a\}$ to node $U\setminus\{a\}$. Indeed, by our previous observations, $T^{\prime}$ is an undirected tree. Furthermore, there is a correspondence between elements of $U$ and the leaves of $T^{\prime}$ (element $b\in U$ corresponds to leaf $\{b\}$ of $T^{\prime}$). To check that $T^{\prime}$ is indeed the laminar tree of $T$, we need to verify that for each bipartition $\{X,Y\}$ in $\mathcal{B}$ there is exactly one edge $e$ of $T^{\prime}$ such that $X=L(T^{\prime}_{1})$ and $Y=L(T^{\prime}_{2})$ for the two connected components $T^{\prime}_{1},T^{\prime}_{2}$ of $T^{\prime}-e$. We say that the edge $e$ implements the bipartition $\{X,Y\}$. Assume that $\{X,Y\}$ is any bipartition from $\mathcal{B}$. First consider the case that $\{X,Y\}$ is the bipartition $\{\{a\},U\setminus\{a\}\}$. Then $\{X,Y\}$ is clearly implemented by the edge $\{\{a\},U\setminus\{a\}\}$ of $T^{\prime}$. Now assume that $\{X,Y\}$ is any other bipartition in $\mathcal{B}$ and assume, without loss of generality, that $a\in Y$. Hence, $X\in\mathcal{S}_{a}$. Let $X^{\prime}$ be the parent of $X$ in $T$. Note that $X$ must be a proper subset of $U\setminus\{a\}$ as by our assumption $\{X,Y\}$ is not the partition $\{\{a\},U\setminus\{a\}$ and hence $X^{\prime}$ cannot be $U$. As we only deleted edge incident to $U$ in our construction of $T^{\prime}$, $\{X,X^{\prime}\}$ is an edge in $T^{\prime}$ which clearly implements the bipartition $\{X,Y\}$. Finally, every bipartition $\{X,Y\}\in\mathcal{B}$ is clearly implemented by only one edge as every inner node of the tree $T^{\prime}$ has degree at least three. Therefore, $T^{\prime}$ is the laminar tree of $(U,\mathcal{B})$.

### v4

@@ 6. Conclusion
We provide transductions for obtaining tree-like graph decompositions such as modular decompositions, cotrees, split decompositions and bi-join decompositions from a graph using $\operatorname{C}{\!\!}\operatorname{MSO}$. This improves upon results of Courcelle [Cou96, Cou06] who gave such transductions for ordered graphs. It is worth mentioning that 22 can be also used to $\operatorname{C}{\!\!}\operatorname{MSO}$-transduce canonical decompositions of other structures such as Tutte’s decomposition of matroids or generally split-decompositions of submodular functions [Cun82] or modular decompositions of $2$-structures [EHR99] or of hypergraphs [HdMMZ22]. As shown by the application given in [Cou06] for transducing Whitney’s isomorphism class of a graph, a line of research is to further investigate which structures can be $\operatorname{C}{\!\!}\operatorname{MSO}$-transduced from a graph or a set system by using the transductions from 2. Also, naturally, the question arises whether counting is necessary or whether $\operatorname{MSO}$ is sufficient to transduce such decompositions.
Furthermore, it is known that if a family of set systems $\mathfrak{S}$ is obtained from a family $\mathfrak{F}$ of $\Sigma$-labeled trees, for some finite alphabet $\Sigma$, by a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\tau$, then for any $\operatorname{C}{\!\!}\operatorname{MSO}$ formula $\varphi$, the set $\{T\in\mathfrak{\mid}\tau(\mathbb{T})\models\varphi\}$ is a recognisable subfamily of $\mathfrak{F}$ [CE12, FMN22]. However, the converse, i.e., whenever a subfamily $\mathfrak{F^{\prime}}$ of $\mathfrak{F}$ is recognisable, then the set $\{\tau(T)\mid T\in\mathfrak{F^{\prime}}\}$ is $\operatorname{C}{\!\!}\operatorname{MSO}$-definable is widely open, and was originally conjectured in [Cou90] for graph classes of small tree-width. Even if Courcelle’s conjecture is now settled [BP16], the case of graphs that are images of $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions, equivalently graph classes of bounded clique-width [CE12], is still open. The $\operatorname{C}{\!\!}\operatorname{MSO}$-definability results shown in this paper, combined with some known special cases, imply that we can push further the known graph classes where recognizability equals $\operatorname{C}{\!\!}\operatorname{MSO}$-definability. In the following subsection we introduce the necessary notation before providing the corollary.
6.1. Recognizability equals definability in restricted classes of graphs of small rank-width
If $G$ is a graph, we denote by $A_{G}$, the adjacency matrix of $G$, a matrix over the binary field whose rows and columns are indexed by $V(G)$ and such that $A_{G}[x,y]=1$ only when $xy$ is an edge. For a graph $G$, we denote by $\rho_{G}:2^{V(G)}\to\mathbb{N}$, the cut-rank function of $G$ where $\rho_{G}(X)$, for $X\subseteq V(G)$, is the rank, over the binary field, of the submatrix of $A_{G}$ whose rows are indexed by $X$ and its columns by $V(G)\setminus X$.
A rank-decomposition of a graph $G$ is a pair $(T,\delta)$ where $T$ is a tree and $\delta:V(G)\to L(T)$ is a bijection from the vertex set of $G$ to the leaves of $T$. The cut-rank of $(T,\delta)$ over $G$ is defined as
$\max\limits_{u\in V(T),X\subseteq N_{T}(u)}\left\{\rho_{G}\left(\bigcup_{v\in X}\delta^{-1}(L(T_{v}^{u}))\right)\right\}.$
The rank-width of a graph $G$ is the maximum cut-rank over all rank-decompositions of $G$. It is known that a graph class $\mathcal{C}$ is included in the image of a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction taking as inputs labeled trees if and only if the rank-width of graphs in $\mathcal{C}$ is bounded by a constant, see for instance [CE12, CGK+25b]. It is also proved in [CGK+25b] that, for every $k$, there are two $\operatorname{C}{\!\!}\operatorname{MSO}$-transductions $\tau_{k}$ and $\mathsf{val}_{k}$ such that $\tau_{k}$ takes as input $\mathbb{G}\sqcup\mathbb{T}$, where $\mathbb{T}$ is a structure describing a rank-decomposition of $G$ of cut-rank at most $k$ and whose leaves are copies of vertices of $G$, and outputs a finitely labeled tree $t$ with $G$ included in $\mathsf{val}_{k}(t)$. Combining this with the following implies that $\operatorname{C}{\!\!}\operatorname{MSO}$-definibility equals recognizability for the following graph classes : distance-hereditary graphs and more generally graphs whose prime graphs are bounded by a constant or have small linear clique-width or have small tree-width or belong to some families of $H$-free graphs such as the once studied in [BDLM05, Rao07].
Theorem 27.
Let $\mathcal{C}$ be a class of graphs for which there is, for every $k$, a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\varphi_{\mathcal{C},k}$, that takes as input a graph $G$ in $\mathcal{C}$ of rank-width at most $k$ and outputs a rank-decomposition of $G$ of width at most $f(k)$, for some computable function $f$. Then, for every positive integer $k$ and for every class of graphs $\mathcal{D}$ whose prime graphs with respect to split decomposition (resp. bi-join decomposition) belong to $\mathcal{C}$, there is a $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\psi_{\mathcal{D},k}$ that takes as input a graph $G$ in $\mathcal{D}$ of rank-width at most $k$ and outputs a rank-decomposition of $G$ of width at most f(k)f(k)(resp. $f(k)+2$).
Let’s recall the Parallel Application Lemma before proceeding to its proof.
[…]
  • the universe of $\bigsqcup_{0<i\leq n}\mathbb{A}_{i}$ is the union of the universes of the $\mathbb{A}_{i}$’s (which are required to be disjoint);
  • for each relation or predicate name from $\Sigma$, its interpretation in $\bigsqcup_{0<i\leq n}\mathbb{A}_{i}$ is the union of its interpretations in each of the $\mathbb{A}_{i}$’s;
[…]
{lemC}
[(Parallel Application Lemma [BGP21])] Let $\tau$ be a $\Sigma$-to-$\Gamma$ $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction. Then there is a $(\Sigma\cup\{\thicksim\})$-to-$(\Gamma\cup\{\thicksim\})$ $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction $\hat{\tau}$ such that, for every sequence $\mathbb{I}_{1},\ldots,\mathbb{I}_{n}$ of $\Sigma$-structures and every sequence $\mathbb{O}_{1},\ldots,\mathbb{O}_{n}$ of $\Gamma$-structures, we have $(\bigsqcup_{0<i\leq n}\mathbb{I}_{i},\bigsqcup_{0<i\leq n}\mathbb{O}_{i})\in\hat{\tau}$ if and only if there exists a permutation $\pi$ of $[n]$ such that $(\mathbb{I}_{i},\mathbb{O}_{\pi(i)})\in\tau$ for all $i\in[n]$.
Proof 6.1 (Proof of 27).
Let $G$ be a connected graph represented by the $\{\mathsf{edge}\}$-structure $\mathbb{G}$. Assume first that we are interested in prime graphs with respect to split decomposition. The transduction is the composition of the following.
  1. (1)
    The first transduction takes as input $\mathbb{G}$ and outputs, thanks to 25, a $\{\mathsf{c}\text{-}\mathsf{edge},\mathsf{t}\text{-}\mathsf{edge}\}$-structure $\mathbb{H}$ which is the canonical split decomposition of $G$.
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
[…]
  5. (5)
    The last transduction takes all rank-decompositions outputted by $\hat{\varphi^{\prime}}$ and fuses any two leaves that are related by a $\mathsf{t}$-edge in $\mathbb{H}$.
It is routine to check that the composition, in the same order, of the above transductions output a rank-decomposition of $G$ as the $\mathsf{t}$-edges form a tree and because each of them yields a correct output when given a correct input. Because the rank-decompositions outputted by $\hat{\varphi^{\prime}}$ have cut-rank at most $f(k)$ on prime subgraphs and of cut-rank $1$ otherwise, it is easy to check that the cut-rank of the outputted rank-decomposition of $G$ has cut-rank at most $f(k)$. If $G$ is not a connected graph, then we apply the above $\operatorname{C}{\!\!}\operatorname{MSO}$-transduction on each connected component thanks again to the Parallel Application Lemma, and then a second non-deterministic transduction adds a new node whose neighbour in the rank-decomposition of each connected component is an arbitrary internal node. The proof for bi-join decompositions works very similar.